The area of a rectangle is 42 m^2, and the length of the rectangle is 5 m more than double the width. Find the dimensions of the rectangle
Answer by Cromlix(4381)

Hi there,
Make the width 'x'
Length = 5 + 2x
Area = Length x Width
42 = (5 + 2x)x
42 = 5x + 2x^2
Sort into ax^2 + bx + c = 0
2x^2 + 5x - 42 = 0
Factorise
(2x - 7)(x + 6)= 0
x + 6 = 0
x = -6 (No answer as -ve)
2x - 7 = 0
2x = 7
x = 7/2 or 3.5
Width = 3.5m
Length = 12m
Hope this helps:-)